OpenStudy (anonymous):

find the equation whose roots are each two more than the roots of x^3 - 4x^2 + 3x - 2 = 0 ?

OpenStudy (shubhamsrg):

let roots be a,b,c now a+b+c=4 ab + bc + ca = 3 abc = 2 the new eqn has roots, a+2,b+2,c+2.. make those substitutions let the new eqn be x^3 + px^2 + qx + r =0 you know, (a+2) + (b+2) + (c+2) = -p (a+2)(b+2) + (b+2)(c+2) + (c+2)(a+2)= q (a+2)(b+2)(c+2) = -r make those comparisons..

OpenStudy (anonymous):

why does a+b+c equal 4?

OpenStudy (shubhamsrg):

sum of roots = -b/a ..rightt ?